# Controllability of the Fisher-Stefan system

### v1

@@ Abstract
Reaction–diffusion system , Fisher–Stefan model , Exact controllability to the trajectories , Free boundary problems for PDEs , Physics-Informed Neural Networks
MSC:
[2020] 35K57 , 80A22 , 93B05 , 93C20 , 35R35 , 68T07
journal: $\alpha\beta\gamma\delta\epsilon\varepsilon\zeta\eta\theta\vartheta\iota\kappa\lambda\mu\nu\xi\pi\varpi\rho\varrho\sigma\varsigma\tau\upsilon\phi\varphi\chi\psi\omega$
[…]
[1]organization=Lab-SIV,addressline=Polydisciplinary Faculty-Ouarzazate, Ibnou Zohr University, city=Ouarzazate, postcode=45000, state=B.P. 638, country=Morocco
[…]
[2]organization=Lab-SIV,addressline=Faculty of Sciences-Agadir, Ibnou Zohr University, city=Agadir, state=B.P. 8106, country=Morocco
[…]
[3]organization=Department of Mathematics,addressline=Faculty of Sciences Semlalia, LMDP, UMMISCO (IRD-UPMC), Cadi Ayyad University, city=Marrakesh, postcode=40000, state=B.P. 2390, country=Morocco
[…]
[4]organization=University Mohammed VI Polytechnic, Vanguard Center,city=Benguerir, country=Morocco
[…]
[5]organization=Department of Applied Mathematics and Statistics. Technical University of Cartagena., city=Campus Muralla del Mar, postcode=30202, state=Cartagena (Murcia), country=Spain
@@ 1.1 Introduction
$\varrho_{t}-\sigma\varrho_{xx}=r\varrho\left(1-\frac{\varrho}{k}\right),$ (1.1)
where $\sigma>0$ is the diffusivity coefficient, $r$ is the growth rate, and $k>0$ is the carrying capacity density. The model (1.1) elucidates the space-time dynamics of a positive population density $\varrho(x,t)$, evolving under the effect of a linear diffusion combined with a logistic source term. See [1, 2, 3, 4] and the references therein.
In this paper, we are interested in a more general situation where the dynamics governed by (1.1) takes place in a right-moving boundary domain $(0,L(t))$. We will assume that the moving boundary evolves according to the following law:
$L^{\prime}(t)=-\mu\varrho_{x}(L(t),t).$ (1.2)
Notably, equation (1.2) resembles the classical Stefan condition [5] used to describe the phase transition of materials. The parameter $\mu$, which corresponds to the inverse Stefan number, serves as a connection between the density gradient at $L(t)$ and the velocity of the boundary. This type of moving-boundary problems have a long history as they have been widely used to simulate a number of physical and industrial processes [6, 7, 8, 9, 10]. In recent years, there has been growing interest in using moving-boundary problems to study biological processes, such as cancer cell invasion, cell movement, and wound healing [11, 12, 13, 14, 15].
@@
$&\varrho_{t}-\sigma\varrho_{xx}=r\varrho\left(1-\frac{\varrho}{k}\right)&&\text{in}\;(0,L(t))\times(0,\infty),\\ &\varrho_{x}(0,t)=0&&\text{in}\;(0,\infty),\\ &\varrho(L(t),t)=0&&\text{in}\;(0,\infty),\\ &L^{\prime}(t)=-\mu\varrho_{x}(L(t),t)&&\text{in}\;(0,\infty),\\ &L(0)=L_{0},\\ &\varrho(\cdot,0)=\varrho_{0}&&\text{in}\;(0,L_{0}),\end{aligned}\right.$ (1.3)
@@
was first proposed by Du and Lin [16], and Du and Guo [17, 18] . In particular, Du and Lin [16] showed that (1.3) has solutions whose population density progresses to a traveling-wave solution with an asymptotic speed that depends on $\mu$. This corresponds to population survival and successful invasion, with $\varrho(x,t)\rightarrow 1$ when $t\rightarrow\infty$. They also proved that (1.3) admits solutions where the population does not settle, so that $\varrho(x,t)\rightarrow 0$ when $t\rightarrow\infty$.
@@
In the context of microbial species, the goal is to apply a control $u(t)$ such that the microbial density $\rho(x,t)$ vanishes across the domain $x\in(0,L(T))$ at a final time horizon $T>0$. Thus, controlling $L(t)$ ensures the complete removal of the microbe. This has important applications in medicine, agriculture, and environmental management. It also falls within the scope of controlling harmful populations or optimizing disinfection. Consequently, we consider (1.3) over a finite time horizon, namely,
@@
$&\varrho_{t}-\sigma\varrho_{xx}=r\varrho\left(1-\frac{\varrho}{k}\right)&&\text{in}\;(0,L(t))\times(0,T),\\ &\varrho_{x}(0,t)=u(t)&&\text{in}\;(0,T),\\ &\varrho(L(t),t)=0&&\text{in}\;(0,T),\\ &L^{\prime}(t)=-\mu\varrho_{x}(L(t),t)&&\text{in}\;(0,T),\\ &L(0)=L_{0},\\ &\varrho(\cdot,0)=\varrho_{0}&&\text{in}\;(0,L_{0}),\end{aligned}\right.$ (1.4)
@@
where $(\varrho_{0},L_{0})$ is the initial state of the system and the control $u(t)$, which is applied at $x=0$, acts as a disinfection flux, impacting the microbial density throughout the domain. In principle, $u(t)$ can be achieved by influencing the boundary flux through $u(t)$ and utilizing diffusion properties to propagate this effect across the entire domain. This type of control might, for instance, represent medical treatment, where antimicrobial substances are administered at a specific point and diffuse through the affected tissue or medium. In agriculture, it could be used to control pathogenic bacteria or fungi in a field or plantation. For more information regarding mathematical models of propagation and control strategies for invasive or harmful species we refer the reader to [19], and, as far as biological and environmental applications of optimal control are concerned, we recommend [20].
@@
Focusing on mathematical controllability properties of free-boundary problem, several recent works [21, 22, 23, 24, 25] have investigated the controllability of one-dimensional parabolic free boundary problems. In [21], a local exact controllability to the trajectories is proved for a system similar to (1.4), with Dirichlet boundary control, but without the nonlinear logistic term. In [22, 23, 24, 25], a local null controllability result for the density is established via a distributed control approach. In [23, 25], the proof is based on fixing the free boundary $L\in C^{1}([0,T])$ and eliminating its dynamics, as well as establishing an observability inequality for the linear heat equation in the non-cylindrical domain $(0,L(t))\times(0,T)$, with an observability constant uniform in $L$. To recover the boundary law, a Schauder fixed-point argument is applied while preserving this law in an appropriate space. In [24], the same local null controllability result is obtained using a different approach, which is based on a non-trivial transformation to a fixed cylindrical domain, as well as a local inversion argument, while in [21] the technique is based on trivial transformation to a fixed cylindrical domain.
@@ 1.2 Main result and reformulation of the problem
$&\varrho_{t}-\varrho_{xx}=\varrho\left(1-\varrho\right)&&\text{in}\;Q_{L},\\ &\varrho_{x}(0,t)=u(t)&&\text{in}\;(0,T),\\ &\varrho(L(t),t)=0&&\text{in}\;(0,T),\\ &L^{\prime}(t)=-\mu\varrho_{x}(L(t),t)&&\text{in}\;(0,T),\\ &L(0)=L_{0},\\ &\varrho(\cdot,0)=\varrho_{0}&&\text{in}\;(0,L_{0}),\end{aligned}\right.$ (1.5)
@@
and we use the definition of smooth trajectories of (1.5):
Definition 1.1.
We call a smooth trajectory of (1.5), any triplet $(\overline{\varrho},\overline{L},\overline{u})$ belonging to
@@
where $\overline{L}\geq L_{0}$ and satisfies:
$&\overline{\varrho}_{t}-\overline{\varrho}_{xx}=\overline{\varrho}\left(1-\overline{\varrho}\right)&&\text{in}\;(0,\overline{L}(t))\times(0,T),\\ &\overline{\varrho}_{x}(0,t)=\overline{u}(t)&&\text{in}\;(0,T),\\ &\overline{\varrho}(\overline{L}(t),t)=0&&\text{in}\;(0,T),\\ &\overline{L}^{\prime}(t)=-\mu\overline{\varrho}_{x}(\overline{L}(t),t)&&\text{in}\;(0,T),\\ &\overline{L}(0)=\overline{L}_{0},\\ &\overline{\varrho}(\cdot,0)=\overline{\varrho}_{0}&&\text{in}\;(0,\overline{L}_{0}).\end{aligned}\right.$
@@
Theorem 1.2.
Let $(\bar{\varrho},\bar{L},\bar{u})$ be a smooth trajectory of (1.5). Then, there exists $\varepsilon>0$ such that for any $L_{0}>0$ and any $\varrho_{0}\in H^{1}(0,L_{0})$ with $\varrho_{0}(L_{0})=0$ satisfying
$|L_{0}-\bar{L}(0)|+\|\varrho_{0}-\bar{\varrho}(\cdot,0)\|_{H^{1}(0,L_{0})}\leq\varepsilon,$
@@
$u(0)=\varrho_{0,x}(0)\quad\mbox{and}\quad L\geq L_{0},$
@@
In our study, we reformulate problem (1.5) as a nonlinear parabolic equation on a fixed cylindrical domain by the change of state:
$\psi(x,t):=\varrho(xL(t),t)\;\mbox{and}\;h(t):=L^{2}(t)\;\mbox{for}\;(x,t)\in Q:=(0,1)\times(0,T).$ (1.6)
@@
{$&h\psi_{t}-\psi_{xx}-\frac{x}{2}h^{\prime}\psi_{x}=h\psi(1-\psi)&&\text{in}\;Q,\\ &\psi_{x}(0,\cdot)=u&&\text{in}\;(0,T),\\ &\psi(1,\cdot)=0&&\text{in}\;(0,T),\\ &h^{\prime}+2\mu\psi_{x}(1,\cdot)=0&&\text{in}\;(0,T),\\ &h(0)=h_{0},\\ &\psi(\cdot,0)=\psi_{0}&&\text{in}\;(0,1),\end{aligned}\right.$ (1.7)
where $h_{0}:=L_{0}^{2}$ and $\psi_{0}(x):=\varrho_{0}(xL_{0})$, for $x$ in $(0,1)$. Consequently, (1.5) is locally exactly controllable to the trajectory $(\bar{\varrho},\bar{L},\bar{u})$ is equivalent that, (1.7) is locally exactly controllable to the trajectory $(\overline{\psi},\overline{h},\overline{u})$ of system (1.7), where $(\bar{\varrho},\bar{L},\bar{u})$ and $(\overline{\psi},\overline{h},\overline{u})$ are linked in compliance with (1.6).
Next, we will transform the local exact controllability to the trajectories of (1.7) into a problem of local null controllability. To accomplish this, we will introduce the change of variable:
@@
where $(\overline{\psi},\overline{h},\overline{u})$ is a trajectory of (1.7) and keep all the terms which are linear with respect to $(z,k)$. The nonlinear system satisfied by the perturbation variables is given by
$&\overline{h}z_{t}-z_{xx}-\frac{x}{2}(\overline{h})^{\prime}z_{x}+\mu x(\overline{\psi})_{x}z_{x}(1,t)+mk+nz=\mathcal{Q}(z,k)&&\text{in}\;Q,\\ &z_{x}(0,t)=\hat{u}(t)&&\text{in}\;(0,T),\\ &z(1,t)=0&&\text{in}\;(0,T),\\ &k^{\prime}(t)+2\mu z_{x}(1,t)=0&&\text{in}\;(0,T),\\ &k(0)=k_{0},\\ &z(\cdot,0)=z_{0}&&\text{in}\;(0,1),\end{aligned}\right.$ (1.8)
where $k^{\prime}$ was replaced by $-2\mu z_{x}(1,t)$ in the first equation of (1.8), $\hat{u}:=u-\overline{u}$, $k_{0}:=h_{0}-\overline{h}(0)$, $z_{0}:=\psi_{0}-\overline{\psi}(\cdot,0)$, the smooth bounded coefficients are given by
$m:=(\overline{\psi})_{t}-\overline{\psi}(1-\overline{\psi}),\quad n:=-\overline{h}(1-2\overline{\psi})$ (1.9)
@@
$\mathcal{Q}(z,k):=-kz_{t}-\mu xz_{x}(1,t)z_{x}-kz^{2}+(1-2\overline{\psi})kz-\overline{h}z^{2}.$ (1.10)
The strategy adopted to control system (1.8) is based on the extension-restriction principle, which transforms a boundary control problem into an interior control problem. This approach involves extending the system to a larger domain where the control acts internally, outside the original domain. Once the solution is controlled in this extended domain, its restriction to the original domain yields a controlled trajectory with Neumann boundary control at $x=0$ and Dirichlet conditions at $x=1$. Hence, we reformulate the local null controllability of (1.8) as local null controllability with interior control by extending the domain $(0,1)$ into $(-1,1)$ with control acts in a region of $(-1,0)$ as shown in Figure 1:
$&\overline{h}z_{t}-z_{xx}-\frac{x}{2}(\overline{h})^{\prime}z_{x}+\mu x(\overline{\psi})_{x}z_{x}(1,\cdot)+mk+nz=\mathcal{Q}(z,k)+\mathds{1}_{\omega}v&&\text{in}\;\tilde{Q},\\ &z(-1,t)=z(1,t)=0&&\text{in}\;(0,T),\\ &k^{\prime}(t)+2\mu z_{x}(1,t)=0&&\text{in}\;(0,T),\\ &k(0)=k_{0},\\ &z(\cdot,0)=z_{0}&&\text{in}\;(-1,1),\end{aligned}\right.$ (1.11)
where $\tilde{Q}:=(-1,1)\times(0,1)$ and $\omega\subset\subset(-1,0)$ is an open subset of $(-1,0)$.
$\omega$11-1- 1001111
Figure 1: The extended domain $(-1,1)$ of $(0,1)$ and the control region $\omega$.
@@
To simplify, we have retained the same notations $z_{0}$, $\overline{\psi}$ and $\overline{h}$ for the new extended data of system (1.11). Specifically, $z_{0}$ has been extended as an element of $H^{1}(-1,1)$, while $\overline{\psi}$ and $\overline{h}$ have been extended as follows:
$(\overline{\psi},\overline{h})\in\left[H^{2,1}(\tilde{Q})\cap W^{1,\infty}(0,T;H^{1}(0,T))\right]\times W^{1,\infty}(0,T),$ (1.12)
which satisfies $\overline{h}\geq L_{0}^{2}$, in order to ensure the regularity required for investigating the Carleman estimate of [21].
Considering all the above, the main Theorem 1.2 will be an immediate consequence of the following theorem:
Theorem 1.3.
@@
$|k_{0}|+\|z_{0}\|_{H^{1}(0,1)}\leq\varepsilon,$
there exists a control $v\in L^{2}(\omega\times(0,T))$ such that the associated state $(z,k)$ of system (1.11) belonging to $H^{2,1}_{0}(\tilde{Q})\times H^{1}(0,T)$ and satisfies:
@@
$H_{0}^{2,1}(\tilde{Q}):=\left\{z\in H^{2,1}(\tilde{Q})\;:\;z(-1,\cdot)=z(1,\cdot)=0\;\mbox{in}\;(0,T)\right\}.$
To achieve this, we reformulate the local null controllability property of system (1.11) as the surjectivity of the mapping $\Lambda:\mathds{X}\longrightarrow\mathds{Y}$, where $\mathds{X}$ and $\mathds{Y}$ are appropriately chosen Hilbert spaces (see Section 3):
$\Lambda(z,k,v):=\left(\Lambda_{1}(z,k,v),\Lambda_{2}(z,k,v),z(\cdot,0),k(0)\right),$ (1.13)
$\Lambda_{1}(z,k,v):=\overline{h}z_{t}-z_{xx}-\frac{x}{2}(\overline{h})^{\prime}z_{x}+\mu x(\overline{\psi})_{x}z_{x}(1,\cdot)+mk+nz-\mathcal{Q}(z,k)-\mathds{1}_{\omega}v,$
$\Lambda_{2}(z,k,v):=k^{\prime}-2\mu z_{x}(\cdot,t).$
More precisely, Theorem 1.3 is equivalent to

### v2

@@ Abstract
Reaction-diffusion system, Fisher-Stefan model, Control to trajectories, Free boundary problems for PDEs, Physics-Informed Neural Networks
@@ 1 Introduction and main results
$\varrho_{t}-\sigma\varrho_{xx}=r\varrho\left(1-\frac{\varrho}{k}\right),$ (1)
where $\sigma>0$ is the diffusivity coefficient, $r$ is the growth rate, and $k>0$ is the carrying capacity density. The model (1) elucidates the space-time dynamics of a positive population density $\varrho(x,t)$, evolving under the effect of a linear diffusion combined with a Logistic source term. See [5, 24, 27, 32] and the references therein.
In this paper, we are interested in a more general situation where the dynamics governed by (1) takes place in a right-moving boundary domain $(0,L(t))$. We will assume that the moving boundary evolves according to the following law:
$L^{\prime}(t)=-\mu\varrho_{x}(L(t),t).$ (2)
Notably, equation (2) resembles the classical Stefan condition [6] used to describe the phase transition of materials. The parameter $\mu$, which corresponds to the inverse Stefan number, serves as a connection between the density gradient at $L(t)$ and the velocity of the boundary. This type of moving-boundary problems have a long history as they have been widely used to simulate a number of physical and industrial processes [3, 7, 30, 31, 4]. In recent years, there has been growing interest in using moving-boundary problems to study biological processes, such as cancer cell invasion, cell movement, and wound healing [13, 21, 25, 37, 39].
@@
$\left\{\begin{aligned} &\varrho_{t}-\sigma\varrho_{xx}=r\varrho\left(1-\frac{\varrho}{k}\right)&&\text{in}\;(0,L(t))\times(0,\infty),\\ &\varrho_{x}(0,t)=0&&\text{in}\;(0,\infty),\\ &\varrho(L(t),t)=0&&\text{in}\;(0,\infty),\\ &L^{\prime}(t)=-\mu\varrho_{x}(L(t),t)&&\text{in}\;(0,\infty),\\ &L(0)=L_{0},\\ &\varrho(\cdot,0)=\varrho_{0}&&\text{in}\;(0,L_{0}),\end{aligned}\right$. (3)
@@
was first proposed by Du and Lin [9], and Du and Guo [10, 11] . In particular, Du and Lin [9] showed that (3) has solutions whose population density progresses to a traveling-wave solution with an asymptotic speed that depends on $\mu$. This corresponds to population survival and successful invasion, with $\varrho(x,t)\rightarrow 1$ when $t\rightarrow\infty$. They also proved that (3) admits solutions where the population does not settle, so that $\varrho(x,t)\rightarrow 0$ when $t\rightarrow\infty$.
@@
In the context of microbial species, the goal is to apply a control $u(t)$ such that the microbial density $\rho(x,t)$ vanishes across the domain $x\in(0,L(T))$ at a final time horizon $T>0$. Thus, controlling $L(t)$ ensures the complete removal of the microbe. This has important applications in medicine, agriculture, and environmental management. It also falls within the scope of controlling harmful populations or optimizing disinfection. Consequently, we consider (3) over a finite time horizon, namely,
@@
$\left\{\begin{aligned} &\varrho_{t}-\sigma\varrho_{xx}=r\varrho\left(1-\frac{\varrho}{k}\right)&&\text{in}\;(0,L(t))\times(0,T),\\ &\varrho_{x}(0,t)=u(t)&&\text{in}\;(0,T),\\ &\varrho(L(t),t)=0&&\text{in}\;(0,T),\\ &L^{\prime}(t)=-\mu\varrho_{x}(L(t),t)&&\text{in}\;(0,T),\\ &L(0)=L_{0},\\ &\varrho(\cdot,0)=\varrho_{0}&&\text{in}\;(0,L_{0}),\end{aligned}\right$. (4)
@@
where $(\varrho_{0},L_{0})$ is the initial state of the system and the control $u(t)$, which is applied at $x=0$, acts as a disinfection flux, impacting the microbial density throughout the domain. In principle, $u(t)$ can be achieved by influencing the boundary flux through $u(t)$ and utilizing diffusion properties to propagate this effect across the entire domain. This type of control might, for instance, represent medical treatment, where antimicrobial substances are administered at a specific point and diffuse through the affected tissue or medium. In agriculture, it could be used to control pathogenic bacteria or fungi in a field or plantation. For more information regarding mathematical models of propagation and control strategies for invasive or harmful species [28], and, as far as biological and environmental applications of optimal control are concerned [38].
@@
Focusing on mathematical controllability properties of free-boundary problem, several recent works [1, 8, 14, 18, 16] have investigated the controllability of one-dimensional parabolic free boundary problems. In [1], a local exact controllability to the trajectories is proved for a system similar to (4), with Dirichlet boundary control, but without the nonlinear Logistic term. In [8, 14, 18, 16], a local null controllability result for the density is established via a distributed control approach. In [14, 16], the proof is based on fixing the free boundary $L\in C^{1}([0,T])$ and eliminating its dynamics, as well as establishing an observability inequality for the linear heat equation in the non-cylindrical domain $(0,L(t))\times(0,T)$, with an observability constant uniform in $L$. To recover the boundary law, a Schauder fixed-point argument is applied while preserving this law in an appropriate space. In [18], the same local null controllability result is obtained using a different approach, which is based on a non-trivial transformation to a fixed cylindrical domain, as well as a local inversion argument, while in [1] the technique is based on trivial transformation to a fixed cylindrical domain.
@@ 1.1 Main result and reformulation of the problem
$\left\{\begin{aligned} &\varrho_{t}-\varrho_{xx}=\varrho\left(1-\varrho\right)&&\text{in}\;Q_{L},\\ &\varrho_{x}(0,t)=u(t)&&\text{in}\;(0,T),\\ &\varrho(L(t),t)=0&&\text{in}\;(0,T),\\ &L^{\prime}(t)=-\mu\varrho_{x}(L(t),t)&&\text{in}\;(0,T),\\ &L(0)=L_{0},\\ &\varrho(\cdot,0)=\varrho_{0}&&\text{in}\;(0,L_{0}),\end{aligned}\right$. (5)
@@
and we use the definition of smooth trajectories of (5):
Definition 1.
We call a smooth trajectory of (5), any triplet $(\overline{\varrho},\overline{L},\overline{u})$ belonging to
@@
where, for all $t\in[0,T]$, $\overline{L}(t)>L_{\star}$ for some fixed constant $L_{\star}>0$ and satisfies:
$\left\{\begin{aligned} &\overline{\varrho}_{t}-\overline{\varrho}_{xx}=\overline{\varrho}\left(1-\overline{\varrho}\right)&&\text{in}\;(0,\overline{L}(t))\times(0,T),\\ &\overline{\varrho}_{x}(0,t)=\overline{u}(t)&&\text{in}\;(0,T),\\ &\overline{\varrho}(\overline{L}(t),t)=0&&\text{in}\;(0,T),\\ &\overline{L}^{\prime}(t)=-\mu\overline{\varrho}_{x}(\overline{L}(t),t)&&\text{in}\;(0,T),\\ &\overline{L}(0)=\overline{L}_{0},\\ &\overline{\varrho}(\cdot,0)=\overline{\varrho}_{0}&&\text{in}\;(0,\overline{L}_{0}).\end{aligned}\right$.
@@
Theorem 1.
Let $(\bar{\varrho},\bar{L},\bar{u})$ be a smooth trajectory of (5). Then, there exists $\varepsilon>0$ such that for any $L_{0}>{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}L_{\star}}$ and any $\varrho_{0}\in H^{1}(0,L_{0})$ with $\varrho_{0}(L_{0})=0$ satisfying
$|L_{0}-\bar{L}(0)|+\|\varrho_{0}-\bar{\varrho}(\cdot,0)\|_{H^{1}(0,L_{0})}<\varepsilon,$
@@
$u(0)=\varrho_{0,x}(0)\quad\mbox{and}\quad L>{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}L_{\star}},$
@@
In our study, we reformulate problem (5) as a nonlinear parabolic equation on a fixed cylindrical domain by the change of state:
$\psi(x,t):=\varrho(xL(t),t)\;\mbox{and}\;h(t):=L^{2}(t)\;\mbox{for}\;(x,t)\in Q:=(0,1)\times(0,T).$ (6)
@@
$\left\{\begin{aligned} &h\psi_{t}-\psi_{xx}-\frac{x}{2}h^{\prime}\psi_{x}=h\psi(1-\psi)&&\text{in}\;Q,\\ &\psi_{x}(0,\cdot)={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sqrt{h}}u&&\text{in}\;(0,T),\\ &\psi(1,\cdot)=0&&\text{in}\;(0,T),\\ &h^{\prime}+2\mu\psi_{x}(1,\cdot)=0&&\text{in}\;(0,T),\\ &h(0)=h_{0},\\ &\psi(\cdot,0)=\psi_{0}&&\text{in}\;(0,1),\end{aligned}\right$. (7)
where $h_{0}:=L_{0}^{2}$ and $\psi_{0}(x):=\varrho_{0}(xL_{0})$, for $x$ in $(0,1)$. Consequently, (5) is locally exactly controllable to the trajectory $(\bar{\varrho},\bar{L},\bar{u})$ is equivalent that, (7) is locally exactly controllable to the trajectory $(\overline{\psi},\overline{h},\overline{u})$ of system (7), where $(\bar{\varrho},\bar{L},\bar{u})$ and $(\overline{\psi},\overline{h},\overline{u})$ are linked in compliance with (6).
Next, we will transform the local exact controllability to the trajectories of (7) into a problem of local null controllability. To accomplish this, we will introduce the change of variable:
@@
where $(\overline{\psi},\overline{h},\overline{u})$ is a trajectory of (7) and keep all the terms which are linear with respect to $(z,k)$. The nonlinear system satisfied by the perturbation variables is given by
$\left\{\begin{aligned} &\overline{h}z_{t}-z_{xx}-\frac{x}{2}(\overline{h})^{\prime}z_{x}+\mu x(\overline{\psi})_{x}z_{x}(1,t)+mk+nz=\mathcal{Q}(z,k)&&\text{in}\;Q,\\ &z_{x}(0,t)=\hat{u}(t)&&\text{in}\;(0,T),\\ &z(1,t)=0&&\text{in}\;(0,T),\\ &k^{\prime}(t)+2\mu z_{x}(1,t)=0&&\text{in}\;(0,T),\\ &k(0)=k_{0},\\ &z(\cdot,0)=z_{0}&&\text{in}\;(0,1),\end{aligned}\right$. (8)
where $k^{\prime}$ was replaced by $2\mu z_{x}(1,t)$ in the first equation of (8), $\hat{u}:={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sqrt{h}}u-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sqrt{\overline{h}}}\overline{u}$, $k_{0}:=h_{0}-\overline{h}(0)$, $z_{0}:=\psi_{0}-\overline{\psi}(\cdot,0)$, the smooth bounded coefficients are given by
$m:=(\overline{\psi})_{t}-\overline{\psi}(1-\overline{\psi}),\quad n:=-\overline{h}(1-2\overline{\psi})$ (9)
@@
$\mathcal{Q}(z,k):=-kz_{t}-\mu xz_{x}(1,t)z_{x}-kz^{2}+(1-2\overline{\psi})kz-\overline{h}z^{2}.$ (10)
The strategy adopted to control system (8) is based on the extension-restriction principle, which transforms a boundary control problem into an interior control problem. This approach involves extending the system to a larger domain where the control acts internally, outside the original domain. Once the solution is controlled in this extended domain, its restriction to the original domain yields a controlled trajectory with Neumann boundary control at $x=0$ and Dirichlet conditions at $x=1$. Hence, we reformulate the local null controllability of (8) as local null controllability with interior control by extending the domain $(0,1)$ into (1,1)(-1,1) with control acts in a region of (1,0)(-1,0) as shown in Figure 1:
$\left\{\begin{aligned} &\overline{h}z_{t}-z_{xx}-\frac{x}{2}(\overline{h})^{\prime}z_{x}+\mu x(\overline{\psi})_{x}z_{x}(1,t)+mk+nz=\mathcal{Q}(z,k)+\mathds{1}_{\omega}v&&\text{in}\;\tilde{Q},\\ &z(-1,t)=z(1,t)=0&&\text{in}\;(0,T),\\ &k^{\prime}(t)+2\mu z_{x}(1,t)=0&&\text{in}\;(0,T),\\ &k(0)=k_{0},\\ &z(\cdot,0)=z_{0}&&\text{in}\;(-1,1),\end{aligned}\right$. (11)
where $\tilde{Q}:=(-1,1)\times(0,1)$ and $\omega\subset\subset(-1,0)$ is an open subset of (1,0)(-1,0).
$\omega$1-1011
Figure 1: The extended domain (1,1)(-1,1) of $(0,1)$ and the control region $\omega$.
@@
To simplify, we have retained the same notations $z_{0}$, $\overline{\psi}$ and $\overline{h}$ for the new extended data of system (11). Specifically, $z_{0}$ has been extended as an element of $H^{1}(-1,1)$, while $\overline{\psi}$ and $\overline{h}$ have been extended as follows:
$(\overline{\psi},\overline{h})\in\left[H^{2,1}(\tilde{Q})\cap W^{1,\infty}(0,T;H^{1}(0,T))\right]\times W^{1,\infty}(0,T),$ (12)
which satisfies $\overline{h}>{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}L_{\star}^{2}}$, in order to ensure the regularity required for investigating the Carleman estimate of [1].
Considering all the above, the main Theorem 1 will be an immediate consequence of the following theorem:
Theorem 2.
@@
$|k_{0}|+\|z_{0}\|_{H^{1}(-1,1)}<\varepsilon,$
there exists a control $v\in L^{2}(\omega\times(0,T))$ such that the associated state $(z,k)$ of system (11) belonging to $H^{2,1}_{0}(\tilde{Q})\times H^{1}(0,T)$ and $\|(z,k)\|_{H^{2,1}_{0}(\tilde{Q})\times H^{1}(0,T)}\leq C(|k_{0}|+\|z_{0}\|_{H^{1}(-1,1)})$ for some constant $C>0$ independent of $\varepsilon$, satisfies:
@@
$H_{0}^{2,1}(\tilde{Q}):=\left\{z\in H^{2,1}(\tilde{Q})\;:\;z(-1,t)=z(1,t)=0\;\mbox{in}\;(0,T)\right\}.$
Remark 2.
By the energy estimate given in Theorem 2 and since $\overline{h}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}>}L_{\star}^{2}$, we can choose $\varepsilon>0$ sufficiently small so that
$|k(t)|<\min_{t\in[0,T]}\overline{h}(t)-L_{\star}^{2}\quad\forall t\in[0,T].$
Consequently, we obtain
$h(t)$ $=k(t)+\overline{h}(t)>L_{\star}^{2}\quad\mbox{and}\quad L(t)=\sqrt{h(t)}>L_{\star}\quad\forall t\in[0,T].$
We now explain why Theorem 1 follows from Theorem 2. Indeed, by the restriction strategy and the trace theorem [29, Theorem 2.1], the control $\hat{u}$ is defined by
$\hat{u}=z_{x}(x,\cdot)\big|_{x=0}\in H^{1/4}(0,T),$
where $(z,k)$ is the controlled state of system (11). The control $u$ appearing in Theorem 1 is then given by
$u=\frac{1}{\sqrt{h}}\Bigl(\hat{u}+\sqrt{\overline{h}}\,\overline{u}\Bigr).$
Since $\overline{u}\in H^{1/4}(0,T)$, $h,\overline{h}\in H^{1}(0,T)$ and satisfy $h(t),\overline{h}(t)>L_{\star}^{2}>0,\;t\in(0,T)$, the functions $1/\sqrt{h}$ and $\sqrt{\overline{h}}$ belong to $H^{1}(0,T)$. Consequently, using the multiplication theorem in Sobolev spaces [2, Theorem 7.4], we deduce that $u\in H^{1/4}(0,T)$.
To achieve this, we reformulate the local null controllability property of system (11) as the surjectivity of the mapping $\Lambda:\mathds{X}\longrightarrow\mathds{Y}$, where $\mathds{X}$ and $\mathds{Y}$ are appropriately chosen Hilbert spaces (see Section 3):
$\Lambda(z,k,v)=\left(\Lambda_{1}(z,k,v),\Lambda_{2}(z,k,v),z(\cdot,0),k(0)\right),$ (13)
$\Lambda_{1}(z,k,v)=\overline{h}z_{t}-z_{xx}-\frac{x}{2}(\overline{h})^{\prime}z_{x}+\mu x(\overline{\psi})_{x}z_{x}(1,\cdot)+mk+nz-\mathcal{Q}(z,k)-\mathds{1}_{\omega}v,$
$\Lambda_{2}(z,k,v)=k^{\prime}-2\mu z_{x}(\cdot,t).$
More precisely, Theorem 2 is equivalent to